# Pooling Liquidity Pools in AMMs

### v1

@@ Abstract
Market fragmentation across multiple Automated Market Makers (AMMs) creates inefficiencies such as costly arbitrage, unnecessarily high slippage and delayed incorporation of new information into prices. These inefficiencies raise trading costs, reduce liquidity provider profits, and degrade overall market efficiency. To address these issues, we propose a modification of the Constant Product Market Maker (CPMM) pricing mechanism, called the Global Market Maker (GMM), which aggregates liquidity information from all AMMs to mitigate these inefficiencies. Through theoretical and numerical analyses, we demonstrate that the GMM enhances profits for both AMMs and traders by eliminating arbitrage opportunities. Additionally, it reduces the profitability of sandwich attacks and minimizes impermanent losses
@@ 1 Introduction
As cryptocurrentcies’ popularity grows, so it does the need to exchange them for other cryptocurrencies and for fiat currencies. Crypto exchanges have emerged to satisfy these needs. A crypto exchange is a platform that allows users to buy, sell, and trade cryptocurrencies. Crypto exchanges can be centralized or decentralized. Centralized EXchanges (CEXs) are owned and operated by a single entity, which holds users’ funds and executes trades. Decentralized EXchanges (DEXs) do not rely on a central authority to hold users’ funds or execute trades. Instead, they use smart contracts running over a public blockchain such as Ethereum, to automate the trading process. DEXs have become increasingly popular reaching a daily trading volume of billions of dollars DEX Tracker - Decentralized Exchanges Trading Volume (2023).
There are two main types of DEXs, namely, order book DEXs and Automated Market Makers (AMMs). Order book DEXs were introduced first but nowadays AMMs are the most common type of DEX. Order book DEXs are similar to traditional order book exchanges. Users place orders to buy or sell assets, and these orders are matched with each other to execute trades. AMMs work by creating liquidity pools, which are essentially large baskets of assets that are held by the exchange. When a user wants to trade an asset, they do not actually trade with another user. Instead, they trade with the liquidity pool. The AMM will then use a predetermined mathematical formula to determine the price of the trade. One of the most widely used AMMs is the Constant Product MM (CPMM). CPMMs use a constant product formula, that ensures that the product of the reserves of two assets in the liquidity pool remains constant. The CPMM’s formula inputs the volume of the trade and the available liquidity and outputs the price of the trade.
@@
Market fragmentation across multiple AMMs presents challenges, the most notable being the reliance on costly arbitrage to equalize prices. Each CPMM independently adjusts swap rates based on its own trading history, discovering asset prices at different paces. If two CPMMs start with the same exchange rate and one executes a trade, their rates will diverge, with larger trades causing greater price differences. These discrepancies create arbitrage opportunities, systematically exploited by arbitrageurs to restore price consistency. A recent study McLaughlin et al. (2023) estimates that arbitrage profits on the Ethereum blockchain alone amount to $170$ million per year, leading to higher costs for traders and reduced profits for liquidity providers, and thus reducing the efficiency of the blockchain ecosystem.
@@ 2 Literature Review
Several alternative mechanisms have been proposed to mitigate MEV and the costs of arbitrage. Canidio & Fritsch (2024) introduce batch trading, a system where orders are executed in batches rather than continuously, thereby reducing arbitrage and sandwich attacks. This mechanism leverages arbitrageurs’ competition to eliminate price manipulation but relies on oracle pricing. Xavier Ferreira & Parkes (2023) study verifiable sequencing rules to prevent MEV and find that while some non-zero miner profits remain, their proposed sequencing rule ensures that user transactions remain unaffected. Canidio & Fritsch (2024) develop a game-theoretic model to differentiate front-running from legitimate trades. Their proposed protocol reduces front-running risk but requires additional message exchanges, which could impact efficiency. Zhou et al. (2021) proposes a modification of AMM algorithms that incorporate optimal on-chain swap routing and arbitrage.
The literature also provides a detailed analysis of the consequences of arbitrage between AMMs. Milionis et al. (2024) identifies arbitrage as a rebalancing cost for LPs in a theoretical model of liquidity provision. Milionis et al. (2023) studies the effect of fees on arbitrage. Mazor & Rottenstreich (2023) explores cross-chain arbitrage in decentralized exchanges, providing an empirical analysis beyond the Ethereum ecosystem. Hansson (2022) examine arbitrage opportunities in Ethereum, identifying \$30 million in cross-exchange and triangular arbitrage profits across 63,168 trading pairs between July 2020 and February 2022.
The profitability of sandwich attacks and more generally the issue of Maximal Extractable Value (MEV) has been widely documented. Daian et al. (2019) introduce the concept of MEV.Canidio & Danos (2024) provide a game-theoretic analysis of sandwich attacks. Qin et al. (2021) estimate that total extractable value over a 32-month period reached \$541 million, highlighting the scale of MEV extraction. Chi et al. (2024) develop methodologies to identify sandwich attacks, finding that \$675 million was extracted before September 2022 and noting that high-volatility tokens are primary MEV targets. Gramlich (2020) provide a comprehensive literature review on MEV, categorizing extraction strategies and countermeasures. Weintraub et al. (2022) analyze MEV extraction in private pools, quantifying the role of flashbots and other sophisticated arbitrage techniques.
More generally, our work is related to the literature on Automated Market Makers (AMMs). Lehar & Parlour (2024) provide a general characterization of AMMs, comparing them to Centralized Exchanges (CEXs). Aoyagi (2023) analyze strategic liquidity provision in AMMs. Cong et al. (2021) analyse the conditions under which CPMM pricing is an optimal liquidity provision mechanism.
@@ 3 Our Benchmark: The Constant Product Market Maker (CPMM)
$\Delta y_{CPMM}(\Delta x_{i};x_{i},y_{i})=\frac{y_{i}}{x_{i}+\Delta x_{i}}\Delta x_{i}=\frac{r_{i}}{1+\frac{\Delta x_{i}}{x_{i}}},$ (2)
@@
where $V_{i,t}\equiv y_{i}+rx_{i}$ is the value of the liquidity holdings of the AMM at the current price, and $V^{0}_{i}\equiv y_{i}^{0}+rx_{i}^{0}$ is the value of the liquidity holdings if the liquidity provider had kept the initial asset allocation. The impermanent losses for the CPMM has a simple expression under the assumption that the initial ratio of reserves of the AMM and the final ratio of reserves are equal to $r_{0}$ and $r$, respectively, and so are the initial and final marginal prices of the AMM:
$\text{IL}_{i}=1-\frac{2}{\sqrt{\frac{r}{r_{0}}}+\sqrt{\frac{r_{0}}{r}}}$
where $r_{0}$ and $r$ are the initial and final price ratios of the token pair, respectively. Figure 2 provides a graph of the impermament losses as a function of the ratio $\frac{r}{r_{0}}$ and the next example illustrates the concept.
Figure 2: The graph plots the impermanent losses as a function of the price increase $\frac{r}{r_{o}}$.
@@
[…]
@@ 4 The Analysis of Global Market Makers
In this section, we study how to twist the design of the CPMM to solve or minimize the problems we discussed in the previous section. Our proposal will be based on a very simple observation. Each AMM has available in blockchains like Ethereum information about the liquidity pools of all other AMMs and this information could be used to determine prices as if the market were not fragmented. We refer these algorithms as global algorithms and we shall focus on global algorithms with certain properties and study their performance.
[…]
@@
We are interested in algorithms that determine the terms of trade usign the information contained in the vector of liquidity reserves or all AMMs. We refer to such algorithms as global. The global algorithm that mimics the behavior of a CPMM algorithm with the aggregate reserves is as follows.
@@
Suppose instead that the a trader sends 10 ETH to AMM 1 when AMM 1’s pool contains 90 ETH and 444,444 UST and AMM 2’s pool 100 ETH and 400000 UST. The trader gets 42,222 UST in return. This is 2,122 ETH more than in the CPMM of Part 1 after arbitrage.
However, the nGMM has two significant drawbacks. The first drawback is immediate: since nGMM sets prices based on the aggregate liquidity, it can lead to complete depletion of one asset’s reserves, forcing the AMM to cease operations. This issue does not arise in the CPMM by design: withdrawing an amount $\Delta y$ of $Y$ arbitrarily close to $y_{i}$ requires supplying an unbounded amount of $X$, causing the per-unit price of $Y$ to diverge to infinity.
The second drawback is more subtle: the nGMM can be exploited if some particular sequences of transactions appear in the market. Specifically, the algorithm may sell an asset at a low price when it is relatively abundant across other AMMs, only to repurchase it at a higher price when it becomes scarce in later trades. As a result, the AMM may suffer a net decrease in its holdings of both assets.
We illustrate these two vulnerabilities with an example.
@@
[frametitle=Toy example - Part 5: Exhausting and Exploiting nGMM] Consider two automated market makers AMM 1 and AMM 2 using nGMM, each with a liquidity pool with 100 ETH and 400000 UST.
  • Exhausting nGMM. Suppose that a trader sends 800,000 UST to AMM 1. According to the nGMM formula, AMM 1 returns 100 ETH to the trader, fully depleting its ETH reserves.
  • Exploiting nGMM. We start again with the initial distribution of reserves of 100 ETH and 400,000 UST in both AMM 1 and AMM 2. Consider the effect of the following transactions:
    • Transaction 1: Trade A sends 10 ETH to AMM 1 and gets 38,095 UST in return. The resulting reserves of AMM 1 are 110 ETH and 361,905 UST. AMM 2 still has 100 ETH and 400,000 UST.
    • Transaction 2: Trade B sends 10 ETH to AMM 2 and gets 34,632 UST in return. The resulting reserves of AMM 2 are then 110 ETH and 365,368 UST. AMM 1 are 110 ETH and 361,905 UST.
@@
    • Transaction 3: Trade C sends 38,095 UST to AMM 1 to get 10.95 ETH. The reserves for AMM 1 would then be 99,05 ETH and 400,000 UST, which is strictly less than its initial reserves (same amount of UST and less of ETH). AMM 2 remains with reserves 110 ETH and 365,368 UST.
@@
    • Transaction 4: Trade D sends 34,632 UST to AMM 2 to get 9.05 ETH in return. The reserves for AMM 2 would then be 100.95 ETH and 400,000 UST, which is strictly greater than its initial reserves (same amount of UST and greater of ETH).
@@
    AMM 2 gains 0.95 ETH that have been obtained from AMM 1. The net variation of trades A, B, C and D is equal to zero. Thus, AMM 2 can still 0.95 ETH from AMM 1 submitting the trades A, B, C and D.
Consequently, the nGMM is not only vulnerable to exploitation, but other AMMs have incentives to do so by submitting zero net-cost orders.
These two vulnerabilities render the nGMM unsuitable for practical implementation. Our approach, therefore, is to explore how much we can adjust the CPMM to approximate the nGMM while avoiding these vulnerabilities.
For this purpose, we adopt a highly conservative approach. We restrict our attention to algorithms that, once implemented by an AMM, say $i$, prevent exhaustion and exploitation for any initial distribution of reserves $(\vec{x}^{0},\vec{y}^{0})$ and any sequence of swaps.
Formally, a sequence of swaps is given by $\{\iota(t),\Delta x_{t},\Delta y_{t}\}_{t=1}^{T}$, where the index $t$ represents the order of transactions, $\iota(t)$ denotes the identity of the AMM handling the swap, and $\Delta x_{t}$ and $\Delta y_{t}$ represent the quantity of $X$ sent to and the quantity of $Y$ returned by AMM $\iota(t)$, respectively. The case in which the AMM receives $Y$ and supplies $X$ is handled by allowing $\Delta x_{t}$ and $\Delta y_{t}$ to take negative values.
We impose only minimal consistency requirements. We say that a sequence of transactions $\{(\iota(t),\Delta x_{t},\Delta y_{t})\}_{t=1}^{T}$, where $\Delta x_{t}>0$, is consistent with the global algorithm of AMM $i$ if at any $t$ such that $i=\iota(t)$, swapping $\Delta x_{t}$ for $\Delta y_{t}$ (or vice versa) is compatible with the algorithm of AMM $i$ and the vector of liquidity reserves updated with the preceding swaps $\{(\iota(s),\Delta x_{s},\Delta y_{s})\}_{s=1}^{t-1}$.
Definition: The global algorithm used by AMM $i\in I$ is non-exhaustible if there exists no distribution of liquidity pools and no sequence of swaps consistent with the algorithm of $i$ that causes AMM $i$ to run out of either asset, i.e., such that $x_{i}=0$ or $y_{i}=0$.
@@
Definition: A global algorithm employed by AMM $i$ is exploitable if there exists an initial distribution of liquidity pools $(x^{0}_{i},y^{0}_{i})$ and a sequence of swaps consistent with AMM $i$’s algorithm such that AMM $i$ ends up with a lower liquidity pool, i.e., $(x_{i},y_{i})\leq(x_{i}^{0},y_{i}^{0})$. [^3] We denote $(x_{i},y_{i})\leq(x_{i}^{\prime},y_{i}^{\prime})$ if $x_{i}\leq x_{i}^{\prime}$ and $y_{i}\leq y_{i}^{\prime}$, with at least one inequality being strict.
[…]
The restriction to algorithms between the CPMM and the nGMM should be interpreted as a minimal departure from existing practices. The CPMM represents the current state of the art, while the nGMM serves as its ideal generalization to mitigate the inefficiencies caused by market fragmentation. Our objective is to approximate the nGMM as closely as possible while deviating as little as possible from the CPMM.
@@
A global algorithm in between the CPMM and the nGMM is not exploitable if and only if it satisfies:
$\Delta y(\Delta x_{i};\vec{x},\vec{y})\leq\Delta y_{CPMM}(\Delta x_{i};x_{i},y_{i}),$ (5)
[…]
@@
By contradiction, suppose a distribution of reserves $(\vec{x}^{0},\vec{y}^{0})$ and a swap that sends $\Delta x_{i}>0$ of $X$ and gets $\Delta y_{i}$ units of $Y$ in return and that violates (5). The definition of CPMM means that the swap $\Delta x_{i}$ decreases the product of reserves, this is,

### v2

@@ Abstract
Market fragmentation across multiple Automated Market Makers (AMMs) creates inefficiencies such as costly arbitrage, unnecessarily high slippage and delayed incorporation of new information into prices. These inefficiencies raise trading costs, reduce liquidity provider profits, and degrade overall market efficiency. To address these issues, we propose a modification of the Constant Product Market Maker (CPMM) pricing mechanism, called the Global Market Maker (GMM), which aggregates liquidity information from all AMMs to mitigate these inefficiencies. Through theoretical and numerical analyses, we demonstrate that the GMM enhances profits for both AMMs and traders by eliminating arbitrage opportunities, and reducing the profitability of sandwich attacks and impermanent losses.
@@ 1 Introduction
As cryptocurrentcies’ popularity grows, so it does the need to exchange them for other cryptocurrencies and for fiat currencies. Crypto exchanges have emerged to satisfy these needs. A crypto exchange is a platform that allows users to buy, sell, and trade cryptocurrencies. Crypto exchanges can be centralized or decentralized. Centralized EXchanges (CEXs) are owned and operated by a single entity, which holds users’ funds and executes trades. Decentralized EXchanges (DEXs) do not rely on a central authority to hold users’ funds or execute trades. Instead, they use smart contracts running over a public blockchain such as Ethereum, to automate the trading process. DEXs have become increasingly popular reaching a daily trading volume of billions of dollars deifprime (2023).
There are two main types of DEXs, namely, order book DEXs and Automated Market Makers (AMMs). Order book DEXs were introduced first but nowadays AMMs are the most common type of DEX. Order book DEXs are similar to traditional order book exchanges. Users place orders to buy or sell assets, and these orders are matched with each other to execute trades. AMMs work by creating liquidity pools, which are essentially large baskets of assets that are held by the exchange. When a user wants to trade an asset, they do not actually trade with another user. Instead, they trade with the liquidity pool. The AMM will then use a predetermined mathematical formula to determine the price of the trade. One of the most widely used AMMs is the Constant Product Market Maker (CPMM). CPMMs use a constant product formula, that ensures that the product of the reserves of two assets in the liquidity pool remains constant. The CPMM’s formula inputs the volume of the trade and the available liquidity and outputs the price of the trade.
@@
Market fragmentation across multiple AMMs presents challenges, the most notable being the reliance on costly arbitrage to equalize prices. Each CPMM independently adjusts swap rates based on its own trading history, discovering asset prices at different paces. If two CPMMs start with the same exchange rate and one executes a trade, their rates will diverge, with larger trades causing greater price differences. These discrepancies create arbitrage opportunities, systematically exploited by arbitrageurs to restore price consistency. A recent study McLaughlin et al. (2023) estimates that arbitrage profits on the Ethereum blockchain alone amount to \$qgv170170 million per year, leading to higher costs for traders and reduced profits for liquidity providers, and thus reducing the efficiency of the blockchain ecosystem.
@@ 2 Literature Review
This paper is related to a relatively new but fast-growing literature on formalizing and studying the optimal design of constant function market makers (CFMMs) in decentralized exchanges. Angeris & Chitra (2020) introduce a general framework for CFMMs, proving their efficiency in aggregating information and setting prices through reserve functions, in the presence of arbitrage. Angeris et al. (2021) provide a detailed analysis of Uniswap’s AMM, showing how liquidity depth, trade size, fees, and arbitrage activity, interact to determine price slippage, efficiency, and risks for liquidity providers. Milionis & Papadopoulos (2022) propose optimal AMM designs tailored to different market goals, such as incentivizing liquidity provision. Aoyagi (2023) develops a game-theoretic model of AMM markets, analyzing how equilibrium and trader surplus depend on AMM parameters and compare these to traditional double auction markets. Similarly, Milionis & Papadopoulos (2023) study equilibrium and efficiency in AMM-based decentralized exchanges, deriving conditions for price efficiency and welfare, comparing AMM with order book models. Evans (2021) linked AMM mechanisms to classical option pricing, revealing that liquidity providers face similar risks to options writers. Lehar & Parlour (2024) provide theoretical and empirical analysis of Uniswap’s AMM, modeling equilibrium pool sizes, liquidity provision, and the role of arbitrageurs in short-run price discovery. They empirically show that AMMs can offer stable liquidity and efficient price alignment with centralized exchanges. Our paper contributes to this literature by proposing a design that incorporates market information to tackle fragmentation and the need for arbitrageurs, while increasing the incentives for liquidity provision and trade.
This paper is also related to the literature that documents and measures inefficiencies that characterize CFMMs, namely arbitrage and impermanent loss for liquidity providers, and MEV operations for traders. Milionis et al. (2024) identifies arbitrage as a rebalancing cost for liquidity providers in a theoretical model of liquidity provision. Milionis et al. (2023) study the effect of fees on arbitrage. Mazor & Rottenstreich (2023) explores cross-chain arbitrage in decentralized exchanges, providing an empirical analysis beyond the Ethereum ecosystem. Hansson (2022) examine arbitrage opportunities in Ethereum, identifying \$30 million in cross-exchange and triangular arbitrage profits across 63,168 trading pairs between July 2020 and February 2022. Daian et al. (2019) introduce the concept of MEV.Canidio & Danos (2024) provide a game-theoretic analysis of sandwich attacks. Qin et al. (2021) estimate that total extractable value over a 32-month period reached \$541 million, highlighting the scale of MEV extraction. Chi et al. (2024) develop methodologies to identify sandwich attacks, finding that \$675 million was extracted before September 2022 and noting that high-volatility tokens are primary MEV targets. Gramlich (2020) provide a comprehensive literature review on MEV, categorizing extraction strategies and countermeasures. Weintraub et al. (2022) analyze MEV extraction in private pools, quantifying the role of flashbots and other sophisticated arbitrage techniques. Moreover, (Evans (2021); Loesch et al. (2021);Cartea, Jaimungal & Wang (2024)) find that AMM liquidity providers frequently incur impermanent loss that is not fully offset by fee income, particularly in volatile markets. Lastly, Cartea et al. (2023) derive the optimal strategy for arbitrageurs. Our design improves on all those fronts and we quantify the improvement using simulations from different AMMs and time periods.
@@
Lastly, our paper contributes to the literature that explores AMM designs to overcome the aforementioned issues. Cartea, Drissi, Sánchez-Betancourt, Siska & Szpruch (2024) propose a general framework that allows liquidity providers to strategically change the quoted marginal price and the impact of an order to that price, maximizing their profits. Their framework allows for the incorporation of information outside the market but do not address potential issues with oracles.  [^1] For an analysis on the manipulation of oracles see Caldarelli (2020). Our model even though less general, allows us to explicitly study ways to incorporate market information, avoiding manipulation, and to study the effects of the proposed mechanism to both liquidity providers and traders. Canidio & Fritsch (2024) introduce batch trading, a system where orders are executed in batches rather than continuously, thereby reducing arbitrage and sandwich attacks. This mechanism leverages arbitrageurs’ competition to eliminate price manipulation but relies on oracle pricing. Xavier Ferreira & Parkes (2023) study verifiable sequencing rules to prevent MEV and find that while some non-zero miner profits remain, their proposed sequencing rule ensures that user transactions remain unaffected. Canidio & Fritsch (2024) develop a game-theoretic model to differentiate front-running from legitimate trades. Their proposed protocol reduces front-running risk but requires additional message exchanges, which could impact efficiency. Zhou et al. (2021) proposes a modification of AMM algorithms that incorporate optimal on-chain swap routing and arbitrage. In contrast our proposed solution requires only information that already exists in the market.
@@ 3 Our Benchmark: The Constant Product Market Maker (CPMM)
$\Delta y_{CPMM}(\Delta x_{i};x_{i},y_{i})=\frac{y_{i}}{x_{i}+\Delta x_{i}}\Delta x_{i}=\frac{r_{i}}{1+\frac{\Delta x_{i}}{x_{i}}}\Delta x_{i},$ (2)
@@
where $V_{i,t}\equiv y_{i}+rx_{i}$ is the value of the liquidity holdings of the AMM at the current price, and $V^{0}_{i}\equiv y_{i}^{0}+rx_{i}^{0}$ is the value of the liquidity holdings if the liquidity provider had kept the initial asset allocation. The impermanent losses for the CPMM has a simple expression under the assumption that the initial ratio of reserves of the AMM and the final ratio of reserves are equal to $r^{\prime}$ and $r$, respectively, and so are the initial and final marginal prices of the AMM:
$\text{IL}_{i}=1-\frac{2}{\sqrt{\frac{r}{r^{\prime}}}+\sqrt{\frac{r^{\prime}}{r}}}$ (4)
where $r^{\prime}$ and $r$ are the initial and final price ratios of the token pair, respectively. Figure 2 provides a graph of the impermament losses as a function of the ratio $\frac{r}{r^{\prime}}$ and the next example illustrates the concept.
Figure 2: The graph plots the impermanent losses as a function of the price increase $\frac{r}{r^{\prime}}$.
@@
$V_{i}^{0}$ $=$ $100\text{ ETH}\times 3,000\text{ UST/ETH}+400,000\text{ UST},$
$V_{i}$ $=$ $115.47\text{ ETH}\times 3,000\text{ UST/ETH}+346,410.16\text{ UST},$
@@ 4 The Analysis of Global Market Makers
In this section, we begin by observing that the issues of the CPMM, as identified in the previous section, could be mitigated if the AMM made better use of the information available on blockchains like Ethereum, particularly the liquidity pools of other AMMs. Arbitrage could be eliminated by computing the AMM price in relation to existing market prices, which, under the CPMM design, can be determined simply by checking reserve balances. The profitability of sandwich attacks depends on slippage, which, in turn, decreases as the AMM’s reserve size increases—at least in the case of the CPMM. If prices were instead computed using aggregate reserves, they would become less sensitive to large orders, thereby reducing slippage and, consequently, the profitability of sandwich attacks. Finally, if prices adjust before a trade order reaches the AMM, they can respond more quickly to market conditions, helping to minimize impermanent losses.
In our formalization, we assume a set $I$ of AMMs, each with liquidity pools described by the vector $(\vec{x},\vec{y})\in\real_{+}^{2I}$, where
@@
We are interested in algorithms that determine the terms of trade using the information contained in the vector of liquidity reserves or all AMMs. We refer to such algorithms as global. An example is the global algorithm that mimics the behavior of a CPMM algorithm with the aggregate reserves.
@@
Suppose instead that the a trader sends 10 ETH to AMM 1 when AMM 1’s pool contains 90 ETH and 444,444 UST and AMM 2’s pool 100 ETH and 400000 UST. The trader gets 42,222 UST in return. This is 2,122 ETH more than in the CPMM of Part 1 after arbitrage. However, the nGMM has a significant drawback: [^4] Another drawback of the nGMM, which we do not discuss extensively, is that it can lead to the complete depletion of one asset’s reserves, forcing the AMM to cease operations. Neither the CPMM nor the algorithm we propose below, the GMM, share this issue. it can be exploited by an strategic party that submits a specific sequences of transactions in the market. This issue is illustrated in the following example.
@@
[frametitle=Toy example - Part 5: Exploiting nGMM] Consider two automated market makers AMM 1 and AMM 2 using nGMM, each with a liquidity pool with 100 ETH and 400000 UST. Consider the effect of the following transactions:
  • Transaction 1: Send 10 ETH to AMM 1 and gets 38,095 UST in return. The resulting reserves of AMM 1 are 110 ETH and 361,905 UST. AMM 2 still has 100 ETH and 400,000 UST.
  • Transaction 2: Send 10 ETH to AMM 2 and gets 34,632 UST in return. The resulting reserves of AMM 2 are then 110 ETH and 365,368 UST. AMM 1 are 110 ETH and 361,905 UST.
@@
  • Transaction 3: Send 38,095 UST to AMM 1 to get 10.95 ETH. The reserves for AMM 1 would then be 99,05 ETH and 400,000 UST, which is strictly less than its initial reserves (same amount of UST and less of ETH). AMM 2 remains with reserves 110 ETH and 365,368 UST.
@@
  • Transaction 4: Send 34,632 UST to AMM 2 to get 9.05 ETH in return. The reserves for AMM 2 would then be 100.95 ETH and 400,000 UST, which is strictly greater than its initial reserves (same amount of UST and greater of ETH).
@@
AMM 2 gains 0.95 ETH that have been obtained from AMM 1. The net variation of trades A, B, C and D is equal to zero.
In this example, AMM 1 suffers a net decline in its holdings because the nGMM sells UST at a low price when it is relatively abundant among the two AMMs, only to repurchase it later at a higher price since
the sale of UST by AMM 2 has made it relatively more scarce. This net decline in AMM 1 holdings correspond to an equal increase in the holdings of AMM 2 because the orders have zero net flow. More generally, exploitation may take a more sophisticated form and involve other agents. The key feature is that the AMM using the nGMM algorithm suffers a net decline in its holdings after a particular sequence of transactions. Since the decline of the AMM is a gain for the other agents affected by the sequence of transactions, at least one of them can do strictly better by submitting this sequence of transactions because their gain must be which unavoidably results into an identical increase in the net position of all the other agents affected by the transactions
Consequently, one only wants to consider global algorithms for which no sequence of trades lead to a net decline in its holdings. To formalise this condition,
This condition captures a minimal requirement for non-exploitability, asserting that no sequence of orders should be able to move AMM $i$ from its original reserve position to a strictly worse one through a cycle of interactions. It is important to note that the definition is intentionally agnostic about the mechanisms or strategies employed by other AMMs or agents in the system. While some algorithms might not even allow a transition from $(\vec{x}^{0},\vec{y}^{0})$ to $(\vec{x},\vec{y})$, the definition requires only the existence of such a transition, which simplifies the analysis by avoiding the need to track intermediate transactions or specify the behavior of other market participants.
@@
Definition: A global algorithm employed by an AMM $i$ is said to be exploitable if there exist an initial distribution of liquidity pools $(\vec{x}^{0},\vec{y}^{0})$ and a final distribution $(\vec{x},\vec{y})$ such that:
  • There exists an order sent to AMM $i$ that, starting from the initial distribution, moves AMM $i$’s reserves to $(x_{i},y_{i})$.
  • There exists another order sent to AMM $i$ that, starting from the final distribution, moves AMM $i$’s reserves to [^5] We denote $(x_{i},y_{i})\leq(x_{i}^{\prime},y_{i}^{\prime})$ if $x_{i}\leq x_{i}^{\prime}$ and $y_{i}\leq y_{i}^{\prime}$, with at least one inequality being strict. $(x_{i}^{\prime},y_{i}^{\prime})\leq(x^{0}_{i},y^{0}_{i})$.
This condition captures a minimal requirement for non-exploitability, asserting that no sequence of orders should be able to move AMM $i$ from its original reserve position to a strictly worse one through a cycle of interactions. It is important to note that the definition is intentionally agnostic about the mechanisms or strategies employed by other AMMs or agents in the system. While some algorithms might not even allow a transition from $(\vec{x}^{0},\vec{y}^{0})$ to $(\vec{x},\vec{y})$, the definition requires only the existence of such a transition, which simplifies the analysis by avoiding the need to track intermediate transactions or specify the behavior of other market participants.
Moreover, the definition is deliberately neutral with respect to the identity of the strategic actor: it does not matter who executes the trades, only that AMM $i$ incurs a loss. In a closed system, any loss in reserves by one participant necessarily implies that the reserves have been absorbed elsewhere. Therefore, the mere existence of a loss implies the presence of a trader or another AMM that benefits—either directly or through redistribution—from the transaction sequence. This justifies the notion of exploitability and underscores the importance of designing AMM algorithms that are robust against such reserve-depleting sequences.
An obvious approach to achieving non-exploitability is to systematically reduce the amount of assets returned to traders. However, this approach is unappealing, as it assumes that doing so incurs no cost to the AMM. In reality, worsening the terms of trade for traders is likely to reduce their participation. Therefore, it makes sense to require non-exploitability while also ensuring that the terms of trade for traders do not deteriorate. Indeed, this aligns with the spirit of our objective: improving the efficiency of the blockchain ecosystem so that both AMMs and traders benefit.
To address this question we start with a few preliminary observations. First, we adopt the point of view that both AMMs and traders improve absorbing part of the arbitrageurs profits. Thus, the benchmark of comparison must have arbitrage. Second, an algorithm that makes arbitrage less profitable must worsen terms of trade. If successful in eliminating arbitrage, this worsening of the terms of trade should affect traders and thus hinder the possibility of improving traders welfare. Except this worsening of the terms of trade in one AMM is compensated with betters terms of trade in another AMM, which suggests that achiving an improvement for trades requires modeling how they route their orders across AMMs. Here we shall assume that traders always trade with the AMM offering the best terms of trade.
Definition: A global algorithm is trade preserving with respect to the CPMM if for any given initial distribution of liquidity $(\vec{x},\vec{y})$, a trader interested in swapping $\Delta x$ units of $X$ after all arbitrage opportunities have been exhausted can obtain better terms of trade when all the AMMs use the global algorithm than when they all use the CPMM algorithm.
At this level of generality, it is necessary to impose some minimum properties on the global algorithms to deduce meaningful characterizations.
Definition: A global algorithm is natural if:
  • (i)
    It offers better terms of trade to traders when its reserve increase while the ratio of reserves and the other AMMs reserves remain constant.
  • (ii)
    It does not display strictly profitable arbitrage opportunities when all AMMs have the same ratio of reserves.
The first property simply says that slipagge decreases with the size of the AMM and the second that arbitrage opportunites can only arise from differences in the ratio of reserves. Both are natural properties that generalize the CPMM.
@@
A natural global algorithm is trade preserving and not exploitable only if for any possible distribution of reserves and swap it does not return the trader more assets than the CPMM.
@@
By contradiction, suppose a distribution of reserves $(\vec{x}^{0},\vec{y}^{0})$ and a swap that $\Delta x_{i}>0$ of $X$ that gets $\Delta y_{i}$ units of $Y$ in return and that return the trader more assets than the CPMM. The definition of CPMM means that the swap $\Delta x_{i}$ decreases the product of reserves, this is,